the two year zero rate is 6 and the three year zero rate is 6.5%. what is the 1-year forward rate after the second year?! assume continuous compounding
Answer #1

We compute the 1-year forward rate after second year to third year with following equation:

€2*2* $2.3*1 = €13*3

where,

r = zero rate

f = forward rate

putting the values:

20.06*2* 082,3*1 = 0.065*3

e23 - ef23 = 70.062 0.06543

1.21531098648973 1.12749685157938

e23 = 1.07788415088463

f = 1.07788415088463 - 1

23 = 0.07788415088463

fig = 7.79%

Thus, 1-year forward rate from 2nd year to 3rd year = 7.79%

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.